# FaithLM: Towards Faithful Explanations for Large Language Models

### v3

@@ Abstract
Large Language Models (LLMs) have become proficient in addressing complex tasks by leveraging their extensive internal knowledge and reasoning capabilities. However, the black-box nature of these models complicates the task of explaining their decision-making processes. While recent advancements demonstrate the potential of leveraging LLMs to self-explain their predictions through natural language (NL) explanations, their explanations may not accurately reflect the LLMs’ decision-making process due to a lack of fidelity optimization on the derived explanations. Measuring the fidelity of NL explanations is a challenging issue, as it is difficult to manipulate the input context to mask the semantics of these explanations. To this end, we introduce FaithLM for explaining the decision of LLMs with NL explanations. Specifically, FaithLM designs a method for evaluating the fidelity of NL explanations by incorporating the contrary explanations to the query process. Moreover, FaithLM conducts an iterative process to improve the fidelity of derived explanations. Experiment results on three datasets from multiple domains demonstrate that FaithLM can significantly improve the fidelity of derived explanations, which also provides a better alignment with the ground-truth explanations.
@@ 1 Introduction
Large language models (LLMs) exhibit remarkable performance in various natural language processing tasks, such as the GPT4 [1], LLaMA [39], and Claude [4]. However, these language models are commonly regarded as intricate black-box systems. The opacity of their internal mechanisms poses a significant challenge when trying to explain their decision-making process. The lack of transparency in LLMs, especially in API-accessed LLM services, inferences contradict the practical requirements of stakeholders and are in opposition to regulatory standards in various domains, such as GDPR [15, 14]. The imperative arises to develop explainability mechanisms for LLMs, particularly for their use in high-stakes applications such as healthcare.
Numerous studies have attempted to enhance the transparency of decision-making processes in LLMs by providing natural language (NL) explanations. Advanced billion-level LLMs, such as GPT4 [1], mainly leverage their reasoning abilities and internal knowledge for complex downstream problems. However, this complexity poses challenges to faithfully explain the underlying reasons behind their decisions with natural language sentences. In this manner, recent advancements are struggling to generate reliable NL explanations for interpreting LLMs [45]. Some work attempt to leverage powerful LLMs [27, 9] with auxiliary information to generate NL sentences or chain-of-thought (CoT) reasoning [22, 34, 8] as the explanations to explain the decision-making process of LLMs. Although existing work emerged that LLMs may possess the ability to self-explain [26, 42], or follow chain-of-thoughts to explain [25], their explanation-generating process usually overlooks the fidelity, a fundamental metric for evaluating the quality of explanations [12, 41]. Thus, the derived NL explanations may not faithfully reflect the decision-making process of LLMs [46, 40]. Measuring the fidelity of NL explanations is a challenging issue, as the NL explanations may provide crucial information beyond the input context. This makes it hard to mask the semantics of NL explanations by manipulating the input tokens.
@@
Table 1: An example of measuring fidelity of NL explanations. The LLM first answers the question in No. Given a faithful explanation “similar poles pull each other away," with its contrary NL explanation, the LLM changes the answer from No to Yes when introduced contrary NL explanation as an extra condition to LLM. This example indicates that the contrary explanation interrupts the LLM’s original prediction process with faithful information, demonstrating that the information in an original explanation is faithful and aligns with LLM’s initial answer.
To overcome this challenge, we propose a method to measure the fidelity of NL explanations. We give an example to convey the motivation in Figure 1. Specifically, the fidelity of an explanation can be measured by leveraging its contrary explanation as extra conditions of the input context, and observing the LLM’s output difference compared with its initial output. Here, a contrary explanation refers to a statement with opposite semantics to the original explanation. By incorporating the contrary explanation to the input context, we can identify an explanation as high fidelity if there is a significant change in the LLM’s output, such as from No to Yes. This change indicates that the crucial information present in the original explanation is substituted with the opposite meaning context in the contrary explanation, where the crucial information is essential to the LLM’s decision-making process. Based on this observation, we propose to extend the applicability of fidelity to the evaluation on NL explanations. This extension follows the integration of contrary explanations to represent the concepts of masking important features in traditional fidelity measurement.
@@
Building upon this approach for measuring the fidelity of NL explanations, we introduce Faithful LLM Explainers (FaithLM) to generate faithful NL explanations for LLMs. Specifically, FaithLM adopts LLMs as explainers to generate the NL explanations and explanation trigger prompts, and iteratively optimizes the derived NL explanations and trigger prompts with the goal of fidelity enhancement. During the iterative process, FaithLM computes the fidelity of each derived explanation and optimized prompt based on our proposed fidelity measurement method, and progressively improves their fidelity through in-context learning. To demonstrate the effectiveness of FaithLM, we conducted several experiments on four different LLMs under three datasets from different domains. FaithLM achieved significantly higher fidelity performance in generating NL explanations and more closely matched the golden explanations compared with state-of-the-art baseline methods. In summary, our contributions can be summarized as follows:
  • Fidelity of NL Explanations: We propose to measure the fidelity of NL explanation by incorporating its contrary explanation into the query and observing LLM output difference.
  • Faithful LLM Explainers: We introduce FaithLM for improving the fidelity of NL explanations, aiming at faithfully explaining the decision-making process of LLMs.
  • Fidelity and Truthfulness: Experimental results on three datasets show that FaithLM can effectively improve the fidelity of NL explanations, which also shows a better alignment with the ground-truth explanations than state-of-the-art baseline methods.
@@ 2.1 Notations and Objectives
We aim to explain the decisions of arbitrary targeted LLMs $f(\cdot)$ with NL explanations in a post-hoc manner. Given an input $\boldsymbol{X}$, the targeted LLMs generate an output $\boldsymbol{Y}=f(\boldsymbol{X})$. Our objective is to produce an NL explanation $\mathcal{E}_{\text{NL}}$ that faithfully explains the reasons behind the prediction of $\boldsymbol{Y}=f(\boldsymbol{X})$. In this work, we employ an LLM as the explainer $\textsl{g}(\cdot)$ to generate the NL explanation $\mathcal{E}_{\text{NL}}=\textsl{g}(\cdot~{}|~{}\boldsymbol{X},\boldsymbol{Y})$. However, the directly generated $\mathcal{E}_{\text{NL}}$ under single-forward passing may not be faithful and accurate, degrading the user’s trust in the prediction made by the targeted LLM. To this end, our work focuses on benefiting the explainer $\textsl{g}(\cdot)$ to generate more faithful NL explanations regarding the decision of $f(\cdot)$ in a post-hoc manner, where $f(\cdot)$ can be either closed-source or open-source LLMs.
2.2 Post-hoc Explanations of LLMs
Recent research on LLM explanations has primarily focused on post-hoc approaches due to the inaccessibility of LLM weights and architectures. One group of studies calculates importance scores for specific tokens [24, 17], and the other line of work generates natural language sentences [21, 7, 9, 28] and CoT reasoning [25, 8] as explanations by leveraging LLMs. Despite the advanced capability of powerful LLMs to generate NL explanations, issues with unreliability and non-fidelity persist. While studies like Chen et al. [6] utilize internal model weights aiming for reliable explanations generation, accessing these weights is challenging with closed-source LLMs. Considering the non-accessibility of LLMs, such as closed-source LLMs, our goal is to faithfully generate NL explanations instead of exploring the internal mechanisms or neurons of LLMs to generate explanations.
2.3 Limitations of Traditional Fidelity on Natural Language Explanations
@@
The fidelity metric measures the fidelity of the given explanation, which is broadly applicable when ground-truth explanations are unavailable. In the NLP scenario, fidelity has been used to evaluate the heatmap-formatted explanations [24, 17], where the heatmap one highlights the important tokens of the input. Specifically, fidelity evaluates the explanation by removing the important tokens from the input $\boldsymbol{X}$ and checking the prediction difference of the targeted LLM. Following the definition of fidelity [29]. Given a sequence of tokens $\boldsymbol{I}=\{t_{1},\cdots,t_{M}\}\subseteq\mathcal{E}_{\text{NL}}$, which is identified as an important component of explanation to the prediction of a targeted LLM $\boldsymbol{Y}=f(\boldsymbol{X})$. Following the traditional fidelity definition, the fidelity can be estimated as follows:
@@
where "$\boldsymbol{X}\setminus\boldsymbol{I}$" denotes the token removal from $\boldsymbol{X}$ in $\boldsymbol{I}$.
If important component $\boldsymbol{T}$ achieves higher fidelity, this demonstrates that $\mathcal{E}_{\text{NL}}$ comprises the crucial tokens that significantly influence the predictions of the targeted LLMs. However, it is challenging to evaluate the fidelity of NL explanations throughout the fidelity defined above, as the critical components in NL explanations may not contain in the input context $\boldsymbol{X}$. For instance, as the example shown in Figure 1, the critical component “similar poles push each other away" in $\mathcal{E}_{\text{NL}}$ does not explicitly appear in the input context. Thus, we cannot simply remove critical components from the question following the traditional definition. Therefore, in this paper, we propose a solution to systematically address this obstacle by removing the critical components from the semantic level instead of the token level.
@@ 3 FaithLM: The Explainer LLM Framework
In this section, we systematically introduce the generative explanation framework, FaithLM, which derives faithful explanations in natural language. The derived explanations are expected to accurately reflect the predictive decision-making process of targeted LLMs with high fidelity. Specifically, we first introduce the Fidelity Evaluator in Section 3.1 to quantitatively assess the fidelity of the NL explanations. Subsequently, FaithLM iteratively optimizes the fidelity of the derived explanations and searches for improved explanation trigger prompts using the proposed Fidelity Evaluator. The optimization frameworks for these processes are detailed in Sections 3.2 and 3.3.
3.1 Fidelity Evaluator for Natural Language Explanations
Due to the limitation of the fidelity metric described in Section 2.3, we introduce the Fidelity Evaluator to directly assess the fidelity of NL explanations, as shown in Figure 1. Specifically, we first extend the original fidelity definition to evaluate the fidelity of NL explanations $\mathcal{E}_{\text{NL}}$. Finally, we introduce the fidelity evaluator that follows the proposed fidelity measurement criteria.
Fidelity of NL Explanations.
@@ 3.1 Fidelity Evaluator for Natural Language Explanations
Figure 1: The framework of Fidelity Evaluator. The evaluator calculates the fidelity scores of the derived explanations based on its contrary explanations.
To assess the fidelity of NL explanations, we extend the traditional fidelity definition to equip it with special constraints regarding NL explanations. To address this challenge, we propose formulating the fidelity of NL explanation as the prediction difference caused by the “contrary explanation". Specifically, the contrary explanation $\neg\mathcal{E}_{\text{NL}}$ is defined as a statement obtained opposite semantics to the given NL explanation $\mathcal{E}_{\text{NL}}$. For instance, if the explanation is "similar poles push each other away," then the contrary explanation would be "similar poles pull each other closer." To estimate the fidelity of $\mathcal{E}_{\text{NL}}$, we use the contrary explanation $\neg\mathcal{E}_{\text{NL}}$ as an extra given condition to the input queries, forming conditional LLM inferences $f(\boldsymbol{X}~{}|~{}\neg\mathcal{E}_{\text{NL}})$. This operation results in different prediction results for targeted LLMs compared to the original predictions, which compels the targeted LLMs to follow contrary information from the explanations of the input queries. In this manner, the fidelity $\mathcal{S}_{\mathcal{E}}$ of $\mathcal{E}_{\text{NL}}$ can be estimated by the prediction difference:
$\mathcal{S}_{\mathcal{E}}:=f(\boldsymbol{X})-f(\boldsymbol{X}~{}|~{}\neg\mathcal{E}_{\text{NL}}).$ (1)
This formulation aligns with the traditional definition of fidelity by observing the prediction difference of the LLM when considering input with and without the component supplied by the explanations.
Fidelity Evaluator.
Unreliable explanations derived from LLMs often present incorrect predictions [45], which means that the relationship between the reliability of the explanations and the correctness of predictions is significantly high. Motivated by the observation, we propose a framework of the Fidelity Evaluator as illustrated in Figure 1. Specifically, we utilize a powerful LLM agent, such as GPT-3.5, to generate the contrary explanation $\neg\mathcal{E}_{\text{NL}}$ regarding the original $\mathcal{E}_{\text{NL}}$. The generation process for $\neg\mathcal{E}_{\text{NL}}$ is guided by the prompt provided in Appendix G. Given the contrary explanation as a condition for the input context, the Fidelity Evaluator computes the fidelity scores according to the difference of output logits $f(\boldsymbol{X})-f(\boldsymbol{X}~{}|~{}\neg\mathcal{E}_{\text{NL}})$. Intuitively, if the targeted LLMs’ output probability changes significantly enough to flip the result (i.e., judgements or choices), it implies that the contrary explanations $\neg\mathcal{E}_{\text{NL}}$ consist of the contrasting key messages against $\mathcal{E}_{\text{NL}}$ that can significantly impact the decision-making process of the targeted LLM. This suggests that the explanation $\mathcal{E}_{\text{NL}}$ contains key components that significantly support the inference process of the targeted LLM.
3.2 FaithLM for Fidelity-enhanced Explanation

### v4

@@ Abstract
Large language models (LLMs) increasingly produce natural language explanations, yet these explanations often lack faithfulness, and they do not reliably reflect the evidence the model uses to decide. We introduce FaithLM, a model-agnostic framework that evaluates and improves the faithfulness of LLM explanations without token masking or task-specific heuristics. FaithLM formalizes explanation faithfulness as an intervention property: a faithful explanation should yield a prediction shift when its content is contradicted. Theoretical analysis shows that the resulting contrary-hint score is a sound and discriminative estimator of faithfulness. Building on this principle, FaithLM iteratively refines both the elicitation prompt and the explanation to maximize the measured score. Experiments on three multi-domain datasets and multiple LLM backbones demonstrate that FaithLM consistently increases faithfulness and produces explanations more aligned with human rationales than strong self-explanation baselines. These findings highlight that intervention-based evaluation, coupled with iterative optimization, provides a principled route toward faithful and reliable LLM explanations. Our source code is available at https://github.com/ynchuang/xLLM.
FaithLM: Towards Faithful Explanations for Large Language Models
Yu-Neng Chuang1thanks: Correspond to Yu-Neng Chuang <ynchuang@rice.edu>, Vladimir Braverman <vova@cs.jhu.edu> and Xia Hu <xia.hu@rice.edu>   Guanchu Wang1   Chia-Yuan Chang2   Ruixiang Tang1 Shaochen Zhong1Fan Yang3Mengnan Du4Xuanting Cai5 Vladimir Braverman56Xia Hu1 Rice University1  Texas A&M University2  Wake Forest University3 New Jersey Institute of Technology4  John Hopkins University5  Google Research6  Meta Platforms, Inc.7
@@ 1 Introduction
Large language models (LLMs) exhibit remarkable performance in various natural language processing tasks, such as the GPT4 Achiam et al. (2023), LLaMA Touvron et al. (2023), and Claude AnthropicAI (2023). However, these language models are commonly regarded as intricate black-box systems. The opacity of their internal mechanisms poses a significant challenge when trying to explain their decision-making process. The lack of transparency in LLMs, especially in API-accessed LLM services, inferences contradict the practical requirements of stakeholders and are in opposition to regulatory standards in various domains, such as GDPR Goodman et al. (2017); Floridi (2019). The imperative arises to develop explainability mechanisms for LLMs, particularly for their use in high-stakes applications such as healthcare. In this work, we focus on “LLM explanation", rather than “LLM reasoning" or “LLM self-refinement", to interpret the model prediction behaviors after providing the final responses. (More illustrations of their discrepancy are in Section 2.2).
Numerous studies have attempted to enhance the transparency of decision-making processes in LLMs by providing natural language (NL) explanations. However, this complexity poses challenges to faithfully explain the underlying explanation behind their decisions with natural language sentences. Recent advancements are struggling to generate reliable NL explanations for interpreting LLMs Ye and Durrett (2022). Some work attempt to leverage powerful LLMs Majumder et al. (2021); Chen et al. (2023c, b) with auxiliary information to generate NL sentences or heatmap of input tokens as model explanations. Although existing work emerged that LLMs may possess the ability to self-explain Madsen et al. (2024), their explanation-generating process usually overlooks the fidelity, a fundamental metric for evaluating the quality of explanations Chuang et al. (2023); Wang et al. (2023). The derived NL explanations may not faithfully reflect “why model generate this answers." Zhao et al. (2023); Turpin et al. (2023) Some work attempts to leverage chain-of-thought (CoT) reasoning steps as the post-hoc model explanation Lyu et al. (2023); Radhakrishnan et al. (2023). However, these reasoning steps are not considered model explanations in the context of post-hoc explanation Tanneru et al. (2024), where post-hoc ones particularly focus on providing faithful explanations for a given generated answer. These CoT steps are produced without a thorough fidelity check (i.e., one that involves masking out the key factors) to ensure they genuinely influence the final answer. These steps are only the intermediate results during the LLM prediction and their fidelity remains unknown. A proper fidelity measurement requires masking the critical features or key messages in the explanation and observing the model’s performance afterward Du et al. (2019), but neither of them are monitored or adopted before claiming CoT reasoning as model explanations. Measuring the fidelity of NL explanations now become a important but challenging issue, as we can monitor and optimization the explanation generation process based on fidelity improvement. The ones may provide crucial information beyond the input context, but the faithful information may appear in semantic levels, making it hard to measure by manipulating the tokens for fidelity measurement.
@@
Table 1: An example of measuring fidelity of NL explanations. The LLM first answers the question in No. Given a faithful explanation “similar poles pull each other away," with its contrary NL explanation, the LLM changes the answer from No to Yes when introduced contrary NL explanation as an extra condition to LLM. This example indicates that the contrary hint interrupts the LLM’s original prediction process with faithful information, demonstrating that the information in an original explanation is faithful and aligns with LLM’s initial answer.
To overcome this challenge, we assess the fidelity of natural language explanations by treating faithfulness as a causal property: if an explanation captures information or hints the model actually uses, then intervening on that information should predictably change the model’s output. Concretely, given an explanation, we construct a contrary hint that expresses the opposite semantics and append it to the input. We then measure fidelity as the resulting prediction shift relative to the original output. A large, directionally consistent change (for example, from No to Yes) indicates that the explanation contained decision-relevant content that the contrary hint displaced. This evaluation mirrors recent work Chen et al. (2025) that tests whether models use and verbalize externally provided hints, and that reads faithfulness from sensitivity under controlled prompt interventions. Our procedure applies the same principle to explanation content, enabling a simple, model-agnostic fidelity measure for free-text rationales
@@
Building upon this new fidelity measurement, we introduce Faithful LLM Explainers (FaithLM) to generate faithful NL explanations for LLMs. Specifically, FaithLM adopts LLMs as explainer to generate the NL explanations and explanation trigger prompts, and iteratively optimizes the derived NL explanations and trigger prompts with the goal of fidelity enhancement. During the iterative process, FaithLM computes the fidelity of each derived explanation and optimized prompt based on our proposed fidelity measurement method, and progressively improves their fidelity through in-context learning. We conducted the experiments on four different LLMs under three datasets. FaithLM achieves significantly higher fidelity in generating NL explanations and more closely matched the golden explanations compared with state-of-the-art baseline methods. Our contributions can be summarized as follows:
  • Intervention-based fidelity. We define fidelity as prediction sensitivity to a contrary hint that semantically opposes the explanation.
  • Faithful LLM Explainers. Our method generates contrary hints, computes fidelity from outputs, and iteratively refines the prompt and explanation to increase faithfulness.
  • Empirical gains. Experimental results show that FaithLM raises measured fidelity and improves alignment with human explanation over baselines.
@@ 2.1 Notations and Objectives
We aim to explain the decisions of arbitrary targeted LLMs $f(\cdot)$ with NL explanations in a post-hoc manner. Given an input $\boldsymbol{X}$, the targeted LLMs generate an output $\boldsymbol{Y}=f(\boldsymbol{X})$. Our objective is to produce an NL explanation $\mathcal{E}_{\text{NL}}$ that faithfully explains the reasons behind the prediction of $\boldsymbol{Y}=f(\boldsymbol{X})$. In this work, we employ an LLM as the explainer $\textsl{g}(\cdot)$ to generate the NL explanation $\mathcal{E}_{\text{NL}}=\textsl{g}(\cdot~|~\boldsymbol{X},\boldsymbol{Y})$. However, the directly generated $\mathcal{E}_{\text{NL}}$ under single-forward passing may not be faithful and accurate, degrading the user’s trust in the prediction made by the targeted LLM. The consistency between $f(\cdot)$ and $\textsl{g}(\cdot)$ is ensured through an iterative optimization process monitored by fidelity scores. To this end, the explainer $\textsl{g}(\cdot)$ to generate more faithful NL explanations regarding the decision of $f(\cdot)$ in a post-hoc manner, where $f(\cdot)$ can be either closed-source or open-source LLMs. FaithLM targets realistic post-hoc interpretability for API-only or black-box LLMs, so we separate the target model $f(\cdot)$ that answers from the explainer $g(\cdot)$ that proposes $E_{NL}$ (and $\lnot E_{NL}$).
2.2 Difference between LLM Explanation and Chain-of-thoughts
Due to the limited accessibility of LLM APIs, recent research on LLM explanations has largely relied on post-hoc explanation approaches (Chen et al., 2023c). However, some studies conflate ’LLM reasoning’ and “LLM self-refinement’ with ’LLM explanations’ when discussing these post hoc LLM explanations, even though these three terms are not identical and with different goals. We illustrate the difference as follows.
LLM reasoning and Chain-of-thoughts
refers to the internal process the model undergoes when it encounters a query or instruction, such as weighting probabilities and generating words step by step plus verification, with the goal of improving performance on reasoning tasks. Some advantages rely on providing chain-of-thought (CoT) reasoning (Lanham et al., 2023; Radhakrishnan et al., 2023; Chen et al., 2023b; Wang et al., 2022) to present the hidden inference steps that the model goes through. These studies show that CoT manuvinakurike2025thoughts can improve reasoning performance, but does not necessarily provide an explanation or even count as explanations of how or why an LLM arrives at its answers (Tanneru et al., 2024), where “good fidelity” in the series of work is typically defined by the alignment between the content of CoT and the final answer (Lyu et al., 2023; Radhakrishnan et al., 2023). Another line of work leverages self-refinement techniques (Lightman et al., 2023; Madaan et al., 2024; Tian et al., 2024), which employ self-reasoning or knowledge supervision as feedback, to iteratively enhance reasoning performance. Although these advancements introduce robust self-feedback loops that effectively boost reasoning accuracy, the “feedback” during optimization is neither necessarily faithful nor equivalent to LLM explanations (Tanneru et al., 2024). Notably, this feedback may be wrong yet still guide LLMs toward a correct reasoning direction. Due to its non-stationary nature, it yields non-faithful outputs when treated as an LLM explanation, which is also very distinct from the goal of “LLM explanation" tasks.
LLM Explanations.
Unlike LLM reasoning and self-refinement, LLM explanation focuses on clarifying why the model provides a particular answer after generating its final decision (Siegel et al., 2024). The concept of fidelity in an LLM explanation (Du et al., 2019; Zhao et al., 2023), which differs from LLM reasoning and self-refinement, refers to whether the model’s prediction would change if the key knowledge provided explanation were removed. If removing the knowledge causes a drastic change in the model’s prediction, we can conclude that the derived explanation is faithful to the LLMs prediction (i.e., the actual reasons that results the predictions). In this work, we focus on LLM explanation, rather than LLM reasoning or self-refinement.
2.3 Limitations of Traditional Fidelity Measurement on NL Explanations
@@
The fidelity metric measures the fidelity of the given explanation, which is broadly applicable when ground-truth explanations are unavailable. In the NLP scenario, fidelity has been used to evaluate the heatmap-formatted explanations Lopardo et al. (2023); Huang et al. (2023), where the heatmap one highlights the important tokens of the input. Specifically, fidelity evaluates the explanation by removing the important tokens from the input $\boldsymbol{X}$ and checking the prediction difference of the targeted LLM. Following the definition of fidelity Miró-Nicolau et al. (2024). Given a sequence of tokens $\boldsymbol{I}=\{t_{1},\cdots,t_{M}\}\subseteq\mathcal{E}_{\text{NL}}$, which is identified as an important component of explanation to the prediction of a targeted LLM $\boldsymbol{Y}=f(\boldsymbol{X})$. Following the traditional fidelity definition, the fidelity can be estimated as:
@@
where $\boldsymbol{X}\setminus\boldsymbol{I}$" denotes token removal from $\boldsymbol{X}$ in $\boldsymbol{I}$.
If important component $\boldsymbol{T}$ achieves higher fidelity, this demonstrates that $\mathcal{E}_{\text{NL}}$ comprises the crucial tokens that significantly influence the predictions of the targeted LLMs. However, it is challenging to evaluate the fidelity of NL explanations throughout the fidelity defined above, as the critical components in NL explanations may not contain in the input context $\boldsymbol{X}$.Some work Lanham et al. (2023) attempts to measure fidelity by modifying the output chain-of-thought (CoT) reasoning to overcome this challenge. However, altering only the output does not guarantee changes in the model’s pre-filling probability and may therefore meet self-consistency, but rather than the definition of fidelity. Thus, we cannot simply remove or modify critical components from the question following the traditional definition. Unlike the previous approaches, we propose a solution to systematically address this obstacle by removing the critical components from the semantic level instead of the token level.
@@ 3 FaithLM: The Explainer LLM Framework
In this section, we systematically introduce the generative explanation framework, FaithLM, which derives faithful explanations in natural language format. The derived explanations are expected to accurately reflect the predictive decision-making process of targeted LLMs with high fidelity after optimizing under FaithLM.
@@
Figure 1: The fidelity evaluation with hint. The evaluator calculates the fidelity scores of the derived explanations based on its contrary hints.
(a) Fidelity-enhanced Explanation
(b) Trigger Prompt Optimization
Figure 2: An overview of FaithLM framework for two differenent optimization objectives. The blue dotted line reveals the trajectory to optimize the NL explanation (Section 3.2), and the red dotted line indicates the trajectory of the explanation trigger prompt optimization (Section 3.3). “Traj. Prompt" denotes the trajectory system prompt shown in Appendix J.
3.1 Fidelity via Contrary Hint Interventions
To reduce penalization from irrelevant tokens, FaithLM computes the probability gap $\Delta p(Y|X)$ only over tokens with non-zero attribution in $f(\cdot)$’s saliency map, thereby isolating fidelity shifts caused by semantically relevant regions.
Contrary Hint Interventions
We evaluate the fidelity of a natural language explanation $\mathcal{E}_{\text{NL}}$ by treating faithfulness as an intervention property on model inference, in line with recent work that tests faithfulness via controlled changes to the information a model conditions on (Chen et al., 2025; Lanham et al., 2023). Let $f(\mathbf{X})$ denote the model prediction on input $\mathbf{X}$. We construct a contrary hint $\neg\mathcal{E}_{\text{NL}}$, defined as a statement whose semantics are opposed to $\mathcal{E}_{\text{NL}}$ (for example, “similar poles pull each other closer” as the contrary of “similar poles push each other away”). We then condition the model on this contrary hint and read out a fidelity score from the intervention-induced prediction shift:
$\mathcal{S}_{\mathcal{E}}:=f(\mathbf{X})-f(\mathbf{X}\mid\neg\mathcal{E}_{\text{NL}}).$
If $\mathcal{E}_{\text{NL}}$ captures information the model actually uses, inserting $\neg\mathcal{E}_{\text{NL}}$ produces a measurable and directionally consistent change in the output distribution, often sufficient to alter the discrete decision. This operationalization supports free-text explanations without token masking or task-specific heuristics. In practice, our Fidelity Evaluator (i) prompts a strong LLM to produce $\neg\mathcal{E}_{\text{NL}}$ from $\mathcal{E}_{\text{NL}}$ under a fixed instruction template and (ii) estimates $\mathcal{S}_{\mathcal{E}}$ from output logits or calibrated probabilities, using the sensitivity to the contrary-hint intervention as the fidelity signal.
Relation to prior faithfulness evaluations. Our evaluation protocol follows the intervention-based paradigm introduced by prior work (Chen et al., 2025) to test whether verbalized rationales genuinely reflect a model’s internal decision process (Chen et al., 2025). That work manipulates the presence and semantics of hints in prompts and measures whether models both use and verbalize those hints, revealing that stated chains of thought are often unfaithful to underlying reasoning. We adopt the same causal-testing philosophy but apply it to explanation content rather than reasoning text. Specifically, the contrary hint serves as a controlled semantic intervention on the information encoded in the explanation, and the resulting change in model prediction quantifies its fidelity. In this formulation, faithfulness is defined as prediction sensitivity: an explanation is faithful if altering its truth value changes the model’s decision. Consequently, our contrary-hint score provides a direct, model-agnostic instantiation of the evaluation approach advocated by Chen et al. (2025), complementing prior intervention tests on CoT text (Lanham et al., 2023) by operating at the level of meaning rather than surface form.
Theorem 1 (Latent-Context Intervention Validity for Faithfulness).
Let $f:\mathcal{X}\!\times\!\mathcal{C}\!\to\!\Delta(\mathcal{Y})$ be a language model mapping an input $X$ and latent context $C$ to a predictive distribution over an output space $\mathcal{Y}$. Let $E_{NL}$ denote a natural-language explanation of $f(X;C)$, and let $\lnot E_{NL}$ denote its contrary hint. Assume that $E_{NL}$ asserts a proposition about a semantic factor $S_{E}=s(X,C)$, where $s(\cdot)$ extracts the decision-relevant concept, which is latent or retrieved, that the explanation verbalizes. Conditioning on $\lnot E_{NL}$ is equivalent to intervening on this factor while holding $(X,C)$ fixed, i.e., $f(X;C\mid\lnot E_{NL})=f(X;C\mid do(S_{E}\!\leftarrow\!\bar{s}))$ for some contradictory value $\bar{s}$, and predictions are invariant to any irrelevant text $R$, so $f(X;C)=f(X\cup R;C)$. Defining $S_{E}(X;C)=D(f(X;C),f(X;C\mid\lnot E_{NL}))$, where $D$ is any strictly proper divergence, we have
$S_{E}(X;C)=0$ $\iff E_{NL}\text{ is non-faithful for }f(X;C),$
$S_{E}(X;C)>0$ $\iff E_{NL}\text{ is faithful for }f(X;C).$
Hence, the contrary-hint score $S_{E}$ constitutes a valid empirical estimator of faithfulness when the decision-relevant content is not contained in the observed input $X$ but arises from latent or retrieved context.
Intuition of the theoretical foundation. Theorem 1 formalizes this causal perspective: if the explanation encodes information that lies on a causal path to the model’s output, then intervening with a contradictory hint necessarily shifts the predictive distribution, yielding a positive contrary-hint score. Conversely, if the explanation is merely correlational or decorative, the intervention leaves the model unchanged and the score remains zero. Thus, the contrary-hint fidelity measure $S_{E}$ causal faithfulness by directly testing whether the model’s output is functionally dependent on the semantics expressed in its own explanation. More discussions and proof are in Appendix A and Corollary 1.
3.2 FaithLM on Fidelity-enhanced Explanation